OpenStudy (anonymous):

Help!!!! Find the solution to the equations. 2x - y = -3 x + y = 0

OpenStudy (anonymous):

anyone:(:(

OpenStudy (tkhunny):

Any particular method we need to employ?

OpenStudy (anonymous):

no, the choices they give me are, 0,0 -1,1 1,-1

OpenStudy (tkhunny):

Surely you have some method to solve... Elimination? Substitution? Other?

OpenStudy (anonymous):

i dont really know, i just started this and have no clue.plus im doing a quiz.

OpenStudy (tkhunny):

I see. That is not particularly encouraging. If you have no way to solve the system, you have only to guess and see if one works. Simply substitute the given values and see if it works. Let's try (0,0) 2(0) - (0) = 0 - 0 = 0 and that is NOT -3 Okay, (0,0) is no good. Try the next one.

OpenStudy (anonymous):

Im sorry. my guess it is 1,-1

OpenStudy (tkhunny):

Prove it. Show your substitution and subsequent algebra.

OpenStudy (anonymous):

2(1)-(-1)=3 would that be it?

OpenStudy (tkhunny):

That's the one equation. They both have to work. If one fails, there is no need to try the second. If one succeeds, we then have to try the second. BOTH must work if it is a solution.